A sample was taken with 15 observations of soil tests from North Carolina rural yards. We would like to determine whether there is reason to believe the nitrate levels differ from the national average. Our sample average shows 8.19 parts per million (ppm) nitrate concentration with a sample standard deviation of 3.95 ppm. Suppose the national average for rural yard is 6.25 ppm.

Conduct the hypothesis test using a 90% confidence level and answer the following questions. For all numbers, please round to the nearest hundredths.

  • What is the standard error?  
  • What are the degrees of freedom for this sample?
  • What is the T-score?
  • Is the p-value less than alpha (type "yes" or "no")?
  • What is your conclusion? If you decide they NC yards differ from the US average, type "differ", otherwise type "none".  
Answer #1

standard error = 3.95/sqrt(15) = 1.0199

degree of freedom = 15 - 1 = 14


Test statistic,
t = (xbar - mu)/(s/sqrt(n))
t = (8.19 - 6.25)/(3.95/sqrt(15))
t = 1.902

P-value Approach
P-value = 0.0779
As P-value < 0.1, reject the null hypothesis.


differ